A rope of length L has mass per unit length y varying according to the function Y(x)= e^x/L . The rope is pulled by a force of constant 1N on a smooth horizontal surface. Find the tension in the rope at x= L/2.

Grade:11

1 Answers

Chetan Mandayam Nayakar
13 years ago

Dear Ayush, 

total mass of rope=∫e(x/L)dx (from x=0 to L) =L(e-1), acceleration =a =force/mass=1/L(e-1), tension at x=L/2 is equal to

 a∫e(x/L)dx (from x= L/2 to L)= √e/(√e +1)

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras